The minute-hand of a clock is 6 cm long. How far does the end of the hand travel in 35 minutes?
Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)


The minute-hWe have to find the and of a clock is 6 cm long. How far does the end of the hand travel in 35 minutes?
We have to find the angle between two positions when it turns 35 deg
6 degrees per minute
35 minutes the angle will be 6*35 = 210 deglength of arc = (theta/360)*2*pi*r
l= (210/360 )* 2*22/7 *6 = 22cm
Or 7 pi cm

Answer by greenestamps(13200)



We don't NEED to find the angle through which the minute hand rotates; that is unnecessary extra work.

The length of the minute hand (the radius of the circle) is 6cm. In 1 hour (60 minutes) the minute hand make one full revolution, traveling a distance of 6*(2pi) = 12pi cm.

35 minutes is 35/60 = 7/12 of an hour; in 35 minutes the distance the tip of the minute hand moves is (7/12)*12pi = 7pi cm.

ANSWER: 7pi cm